OpenStudy (ell):

Help please! :) Solve each equation by completing the square: a^2-2a-8=0

OpenStudy (anonymous):

a square binomial \[(x+n)^2= x^2+2nx+n^2\] start by moving the 8 to the other side

OpenStudy (ell):

mmkay :) thanx

OpenStudy (anonymous):

a^2-2a = 8 then on the left side, we need to complete the square binomial so we need to determine the 3rd term since the 2nd term is 2a we can conclude n =1 this means that the 3rd term n^2 is 1 at this point, you add 1 to both sides

OpenStudy (ell):

so, add 1 to 8 and that equals 9..... a^2-2a=9

OpenStudy (ell):

@waterineyes could you possibly help?

OpenStudy (anonymous):

add to both sides, not just one

OpenStudy (anonymous):

i moved the -8 to the right side by adding 8 to BOTH !!!! sides when you're adding 1 to complete the square, you need to add it to BOTH!!! sides

OpenStudy (ell):

haha:) mmkay thanx

OpenStudy (anonymous):

so add 1 to both sides to a^2-2a = 8 and what do you get?

OpenStudy (ell):

1a^2-3a=9 ?

OpenStudy (anonymous):

... you cant add 1 to the variables

OpenStudy (anonymous):

you just add it to the left side, not to the terms....

OpenStudy (anonymous):

Where are you facing problem?

OpenStudy (ell):

the problem is a^2-2a-8=0

OpenStudy (anonymous):

Okay.. Here just tell me what is the coefficient of a term??

OpenStudy (ell):

its okay, i got it. Thanks though!!